8. In a lottery 200 tickets are sold. Tickets are either red or blue. The winning ticket is chosen at random. a) What is the probability that the winning ticket will be green? b) 79 children and 90 women buy one ticket each. Men buy the rest of the tickets.

Find the probability that a man buys the winning ticket.


1
Expert's answer
2021-05-27T17:44:29-0400

a) The probability that the winning ticket will be green is:

P(green) = 0/200 = 0.

b) The number of men is: 200 - 79 - 90 = 31.

The probability that a man buys the winning ticket is:

P(man win) = 31/200 = 0.155.